21) Tinas Corp. provided the following partial balance sheet: TINAS CORPS Balance Sheet (partial) Stockholders Equity: $6,
Answer #1

Unrealized gain or loss arising on investment available for sale is recorded in Statement of Comprehensive Income for the year.

1) Option C: Unrealized gain on available for sale securities would increase the comprehensive income.

Comprehensive income represents both Unrealised gain/loss along with tax effect.

2) Option D Loss of $36,400.

Total loss = 56000; Tax rate = 35%

In this unrealized loss arose, therefore, tax-saving occurs which is 56000*35% = 19,600.

Net amount in comprehensive income = 56,000 - 19,600 = $36,400